# An Information Theory Treatment of Animal Movement Tracks

### v2

@@ Abstract
The two-dimensional track of an animal on a landscape has progressed over the past three decades from hourly to second-by-second recordings of locations. Track segmentation methods for analyzing the behavioral information in such relocation data has lagged somewhat behind, with scales of analysis currently at the sub-hourly to minute level. A new approach is needed to bring segmentation analysis down to a second-by-second level. Here, such an approach is presented that rests heavily on concepts from Shannon’s Information Theory. In this paper, we first briefly review and update concepts relating to movement path segmentation. We then discuss how cluster analysis can be used to organize the smallest viable statistical movement elements (StaMEs), which are $\mu$ steps long, and to code the next level of movement elements called “words” that are $m\mu$ steps long. Centroids of these word clusters are identified as canonical activity modes (CAMs). Unlike current segmentation schemes, the approach presented here allows us to provide entropy measures for movement paths, compute the coding efficiencies of derived StaMEs and CAMs, and assess error rates in the allocation of strings of $m$ StaMEs to CAM types. In addition our approach allows us to employ the Jensen-Shannon divergence measure to assess and compare the best choices for the various parameters (number of steps in a StaME, number of StaME types, number of StaMEs in a word, number of CAM types), as well as the best clustering methods for generating segments that can then be used to interpret and predict sequences of higher order segments. The theory presented here provides another tool in our toolbox for dealing with the effects of global change on the movement and redistribution of animals across altered landscapes.
@@ 1 Introduction
These movement elements themselves cannot be identified by their mechanical signature (i.e., fundamental movement elements in the sense described in [2]) because such information is not coded in relocation data alone, but requires either video, accelerometer or other body-parts movement data for identification [3]. Rather they can be identified statistically in the relocation data using a recent approach described in Getz et al. [4]. This approach, though, requires relatively high frequency relocation data that has an inter-point interval of, at most, seconds rather than minutes (i.e. frequencies in excess of $0.1$ Hz).
Decoding a movement track to read the story it encodes involves a process called path segmentation [5]. This process relies on time series segmentation methods such as behavioral change point analysis (BCPA) [6, 7, 8, 9, 10, 11] and hidden Markov methods (HMM) [12, 13, 14, 15, 16]. The latter, as mentioned by Gundermann et al. [10], may not be the best way to identify rare events, such as “parturition, migration initiation, and juvenile dispersal.”
@@ 2 Segmentation Hierarchy
A hierarchical scheme for the segmentation of tracks has been discussed in considerable depth elsewhere [18, 25, 4]. An updated summary is provided in Table 1. Originally, Getz and Saltz [2] proposed that, at its most basic level, a movement track $\mathcal{T}$ can be conceptualized as a sequence of fundamental movement elements (FuMEs; referred to as FMEs in [2]). These elements when strung together either homogeneously (or one type), or in characteristic mixes of several types, produce canonical activity mode (CAMs) segments. Getz and Saltz [2] pointed out that “with most current data, it is not possible to construct distributions of building block movement elements in terms of FuMEs” (they used the acronym FME in [2]): essentially, as previously mentioned, FuMEs are unobservable in relocation data alone and companion video or accelerometer data are likely needed to identify the start of each new FuME. This subsequently led to the concept of a set $\mathcal{S}$ of statistical movement elements (StaMEs, formerly referred to as a metaFuMEs [18] to be used in place of FuMEs as a set of lowest level movement building block elements or coding symbols. In the formulation here, the set $\mathcal{S}$ is regarded as an underlying set of symbols that can be used to encode the movement track relocation data time series $\mathcal{T}^{\rm loc}$ as a StaMEs time series $\mathcal{T}^{\sigma}$ (Table 2). This symbolic representation then allows an information theory toolbox to be used to analyze the information content and, ultimately, meaning of animal movement tracks [30].
@@ 3 Two Level Track Segmentation
    In the same way that ${\mathcal{T}}^{\rm loc}$ can be parsed into $N^{\rm seg}$ segments each consisting of $\mu$ steps, we can parse ${\mathcal{T}}^{\rm loc}$ into $N^{\rm wd}$ each consisting of $m$ consecutive segments providing us with a next level of segmentation in which segments contain $m\mu$ consecutive steps. In this case we obtain a time series of $N^{\rm wd}=\lfloor T/(m\mu)\rfloor$ words
@@ 4 Information Entropy of Movement Track Segments
Information theory, as developed by Claude Shannon [22, 23] in the late 1940’s, posits a set of $n$ symbols ${\mathcal{S}}$ (Eq 3.3) that when strung together sequentially encodes information at a basic rate determined by the probability $p^{\sigma}_{i}$ of encountering symbol $\sigma_{i}$ at randomly selected locations along an encoded string. Under these circumstances the information coding rate of the symbols in $\mathcal{S}$ is determined its $\log_{2}$ entropy $H\left({\mathbf{p}}^{\sigma}\right)$ [37]
@@
where the subscript 0 is used to denote the fact that this compution assumes no autocorrelation among consecutive words. If consecutive words are autocorrelated, however, then the coding rate must account for the fact that the conditional probability $p^{\omega}_{\ell^{\prime}|\ell}$ of the next word being ${\omega}_{\ell^{\prime}}$ when the previous word is $\omega_{\ell}$. In this case, if $p^{\omega}_{\ell\ell^{\prime}}$ is the joint probability for the co-occurrence of the word sequence $\omega_{\ell}$ and $\omega_{\ell^{\prime}}$ independent of the order in which they occur then information content is modified to
@@ 5 Information Content and Method Comparisons
First, we expect that the number of StaMEs groups that can be clearly separated increases with the segment size parameter $\mu$. If we ignore orientation, then for $\mu=2$ only the size (i.e., step length) of StaMEs can vary. If $\mu=3$, then both size and turning angle can vary: apart from size, some StaMEs may have highly acute (almost reversed) turning angles, others closer to right angles, and still others with more obtuse or even rather small turning angles. The latter is characteristic of what is called persistent movement. As $\mu$ increase beyond 3, StaME shapes can become more varied with saw-tooth like elements, circular elements or combinations thereof, and absolute size (i.e., step-size/velocity averaged across all relocation points in the segment) coming into the mix. The more CAMs that are identified, the greater the potential coding rate since the maximum coding rate of a set of $k$ rectified CAMs occurs when $p^{\kappa^{\star}}_{c}=1/k$, $c=1,\cdots,k$ (Eq 3.8), and is given by
@@
Beyond using the entropy measures $H^{\kappa^{\star}_{a}}$ (Eq 4.4) and misassignment error rates $E^{\mathcal{\kappa}_{a}}$ (Eq 3.16) of different functors $a$ to compare various approach and identify those with high coding and low error rates, we can compare the CAM distributions ${\bf p}^{\kappa^{\star}}$ in sets $\mathcal{K}^{\star}_{a}$ that have the same number $k$ of CAMs using the information theoretic Kullback-Leibler and Jensen-Shannon divergence measures [37]. The Kullback-Leibler divergence measure allows us to compute the “distances” (divergences) between two distributions. Specifically, if $\mathbf{p}^{1}=\left(p^{1}_{1},\cdots,p^{1}_{k}\right)$ and $\mathbf{p}^{2}=\left(p^{2}_{1},\cdots,p^{2}_{k}\right)$ are any two discrete distributions with $k$ bins then the Kullback-Leibler measure of the divergence of distribution $\mathbf{p}^{2}$ from distribution $\mathbf{p}^{1}$ is
@@ 6 Discussion
When auxilliary data such as accelerometer, heart rate, skin temperature and barometric readings are collected from an individual along with GPS location data [3, 51, 17], then such data can be incorporated into the formulation to fine tune the informational aspects of movement behavior. The most obvious way to do this is to condition the occurrence and autocorrelations associated with probabilities and sequencing of StaMEs within words, and CAMs within BAMs and to incorporate these changes into the various informational and divergence measures presented here.
@@
The informational approach presented here to segmenting movement tracks of animals, as a way of identifying various factors that influence the structure of animals movement path, provides a rigorous basis for comparing movement tracks of individuals as a function internal, external and idiosyncratic factors that ultimately determine the movement behavior and trajectory of individuals. The theory and methods articulated provide for the first time a way to rigorously address such questions as “Does the information content of an individual increase as it learns to navigate its landscape?” or “Is the information content of an individual reduced when it is infected with parasites?”
At the cusp of the “big data” information age in movement ecology [1], it is only fitting that we have a way of coding movement tracks and assessing and comparing the efficacy of these codes across methods and among movement tracks. This paper provides us with a way forward to achieve this.
@@ Acknowledgments
I would like to thank Varun Sethi for constructive discussions regarding the development of the approach taken here and for taking the time to provide comments on this manuscript, as well as Richard Salter and Orr Spiegel for discussions over the past year regarding animal track segmentation. In addition, and in approximate chronological order, I have benefited over the past 20 years from the many discussions I have had with the following students and colleagues who have influenced my thinking within the field of spatial and movement ecology: Norman Owen-Smith, Jessica Redfern, Peter Baxter, Chris Wilmers, Paul Cross, Jamie Lloyd-Smith, George Wittemyer, Sadie Ryan, Shirli Bar-David, David Saltz, Ran Nathan, Craig Tambling, Leo Polansky, Andy Lyons, Wendy Turner, Steve Bellan, Alex Boettiger, Richard Salter, Werner Kiliam, Royi Zidon, Scott Fortmann-Roe, Miri Tsalyuk, Briana Abrams, Dana Seidel, Eric Dougherty, Colin Carlson, Krti Tallam, Jason Blackburn, Wayne Linklater, Yen-Hua Huang and Ludovica Luisa Vassat,

### v4

@@ Abstract
Position recordings of the two-dimensional tracks of animals moving over landscapes has progressed over the past three decades from hourly to second-by-second locations. Track segmentation methods for analyzing the behavioral information in such relocation data has lagged somewhat behind, with scales of analysis currently at the sub-hourly to minute level. A new approach is needed to bring segmentation analysis down to a second-by-second level. Here, a fine-scale approach is presented that rests heavily on concepts from Shannon’s Information Theory. In this paper, we first briefly review and update concepts relating to movement path segmentation. We then discuss how cluster analysis can be used to organize the smallest viable statistical movement elements (StaMEs), which are $\mu$ steps long, and to code the next level of movement elements called “words” that are $m\mu$ steps long. Centroids of these word clusters are identified as canonical activity modes (CAMs). Unlike current behavioral change point analysis and hidden Markov model segmentation schemes, the approach presented here allows us to provide entropy measures for movement paths, compute the coding efficiencies of derived StaMEs and CAMs, and to assess error rates in the allocation of strings of $m$ StaMEs to CAM types. In addition our approach allows us to employ the Jensen-Shannon divergence measure to assess and compare the best choices for the various parameters (number of steps in a StaME, number of StaME types, number of StaMEs in a word, number of CAM types), as well as the best clustering methods for generating segments that can then be used to interpret and predict sequences of higher order segments. The theory presented here provides another tool in our toolbox for dealing with the effects of global change on the movement and redistribution of animals across altered landscapes.
@@ 1 Introduction
These movement elements themselves cannot be identified by their mechanical signature (i.e., fundamental movement elements in the sense described in [2]) because such information is not coded in relocation data alone, but requires either video, accelerometer or other body-parts movement data for identification [3]. Rather they can be replaced by statistically derived movement elements computed from a fixed number of relocation data steps, using a recent approach described in Getz et al. [4]. This approach permits a coding theory of the relocation data to be formulated, as developed here, though the approach requires relatively high frequency relocation data that has an inter-point interval of, at most, seconds rather than minutes (i.e. frequencies in excess of $0.1$ Hz).
Decoding a movement track to read the story it encodes involves a process called path segmentation [5]. This process relies on time series segmentation methods such as behavioral change point analysis (BCPA) [6, 7, 8, 9, 10, 11] and hidden Markov methods (HMM) [12, 13, 14, 15, 16] that produce variable length segments. Typically, such segments are much longer than the statistical elements discussed in Getz et al. [4]; and we note that if a code is not based on fixed-length segments, then ambiguity arises regarding the number of elements used to code any given segment of track. Additionally, BCPA and HMM methods, as mentioned by Gundermann et al. [10], may not be the best way to identify rare events, such as “parturition, migration initiation, and juvenile dispersal.”
@@ 2 Segmentation Hierarchy
A hierarchical scheme for the segmentation of tracks has been discussed in considerable depth elsewhere [18, 25, 4]. An updated summary is provided in Table 1. Originally, Getz and Saltz [2] proposed that, at its most basic level, a movement track $\mathcal{T}$ can be conceptualized as a sequence of fundamental movement elements (FuMEs; referred to as FMEs in [2]). These elements when strung together either homogeneously (or one type), or in characteristic mixes of several types, produce canonical activity mode (CAMs) segments. Getz and Saltz [2] pointed out that “with most current data, it is not possible to construct distributions of building block movement elements in terms of FuMEs” (they used the acronym FME in [2]): essentially, as previously mentioned, FuMEs are unobservable in relocation data alone and companion video or accelerometer data are likely needed to identify the start of each new FuME. This subsequently led to the concept of a set $\mathcal{S}$ of statistical movement elements (StaMEs), formerly referred to as a metaFuMEs [18] to be used in place of FuMEs as a set of lowest level movement building block elements or coding symbols. In the formulation here, the set $\mathcal{S}$ is regarded as an underlying set of symbols that can be used to encode the movement track relocation data time series $\mathcal{T}^{\rm loc}$ as a StaMEs time series $\mathcal{T}^{\sigma}$ (Table 2). This symbolic representation then allows an information theory toolbox to be used to analyze the information content and, ultimately, meaning of animal movement tracks [30].
@@ 3 Two Level Track Segmentation
    In the same way that ${\mathcal{T}}^{\rm loc}$ can be parsed into $N^{\rm seg}$ segments each consisting of $\mu$ steps, we can parse ${\mathcal{T}}^{\rm loc}$ into $N^{\rm wd}$ words each consisting of $m$ consecutive segments providing us with a next level of segmentation in which segments contain $m\mu$ consecutive steps. In this case we obtain a time series of $N^{\rm wd}=\lfloor T/(m\mu)\rfloor$ words
@@ 4 Information Entropy of Movement Track Segments
Information theory, as developed by Claude Shannon [22, 23] in the late 1940’s, posits a set of $n$ symbols ${\mathcal{S}}$ (Eq 3.3) that when strung together sequentially encodes information at a basic rate determined by the probability $p^{\sigma}_{i}$ of encountering symbol $\sigma_{i}$ at randomly selected locations along an encoded string. Under these circumstances the information coding rate of the symbols in $\mathcal{S}$ is determined by its $\log_{2}$ entropy $H\left({\mathbf{p}}^{\sigma}\right)$ [37]
@@
where the subscript 0 is used to denote the fact that this compution assumes no autocorrelation among consecutive words. If consecutive words are autocorrelated, however, then the coding rate must account for the conditional probability $p^{\omega}_{\ell^{\prime}|\ell}$ of the next word being ${\omega}_{\ell^{\prime}}$ when the previous word is $\omega_{\ell}$. In this case, if $p^{\omega}_{\ell\ell^{\prime}}$ is the joint probability for the co-occurrence of the word sequence $\omega_{\ell}$ and $\omega_{\ell^{\prime}}$ independent of the order in which they occur then the information content is modified to
@@ 5 Information Content and Method Comparisons
First, we expect that the number of StaMEs groups that can be clearly separated increases with the segment size parameter $\mu$. If we ignore orientation, then for $\mu=2$ only the size (i.e., step length) of StaMEs can vary. If $\mu=3$, then both size and turning angle can vary: apart from size, some StaMEs may have highly acute (almost reversed) turning angles, others closer to right angles, and still others with more obtuse or even rather small turning angles. The latter is a characteristic of what is called persistent movement. As $\mu$ increases beyond 3, StaME shapes can become more varied with saw-tooth like elements, circular elements or combinations thereof, and absolute size (i.e., step-size/velocity averaged across all relocation points in the segment) coming into the mix. The more CAMs that are identified, the greater the potential coding rate since the maximum coding rate of a set of $k$ rectified CAMs occurs when $p^{\kappa^{\star}}_{c}=1/k$, $c=1,\cdots,k$ (Eq 3.8), and is given by
@@
Beyond using the entropy measures $H^{\kappa^{\star}_{a}}$ (Eq 4.4) and misassignment error rates $E^{\mathcal{\kappa}_{a}}$ (Eq 3.16) of different functors $a$ to compare various approaches and identify those with high coding and low error rates, we can compare the CAM distributions ${\bf p}^{\kappa^{\star}}$ in sets $\mathcal{K}^{\star}_{a}$ that have the same number $k$ of CAMs using the information theoretic Kullback-Leibler and Jensen-Shannon divergence measures [37]. The Kullback-Leibler divergence measure allows us to compute the “distances” (divergences) between two distributions. Specifically, if $\mathbf{p}^{1}=\left(p^{1}_{1},\cdots,p^{1}_{k}\right)$ and $\mathbf{p}^{2}=\left(p^{2}_{1},\cdots,p^{2}_{k}\right)$ are any two discrete distributions with $k$ bins then the Kullback-Leibler measure of the divergence of distribution $\mathbf{p}^{2}$ from distribution $\mathbf{p}^{1}$ is
@@ 6 Discussion
When auxiliary data such as accelerometer, heart rate, skin temperature and barometric readings are collected from an individual along with GPS location data [3, 51, 17], then such data can be incorporated into the formulation to fine tune the informational aspects of movement behavior. The most obvious way to do this is to condition the occurrence and autocorrelations associated with probabilities and sequencing of StaMEs within words, and CAMs within BAMs and to incorporate these changes into the various informational and divergence measures presented here.
@@
The informational approach presented here to segmenting movement tracks of animals, as a way of identifying various factors that influence the structure of animals movement path, provides a rigorous basis for comparing movement tracks of individuals as a function of internal, external and idiosyncratic factors that ultimately determine the movement behavior and trajectory of individuals. The theory and methods articulated provide for the first time a way to rigorously address such questions as “Does the information content of an individual increase as it learns to navigate its landscape?” or “Is the information content of an individual reduced when it is infected with parasites?”
At the cusp of the “big data” information age in movement ecology [1], it is only fitting that we have a way of coding movement tracks and assessing and comparing the efficacy of these codes across different parameter sets and clustering algorithms. Further, as other kinds of data derived from various types of inertial sensors (e.g., gyroscopes, magnetometers, accelerometers) are collected, such data can be used to supplement location data in characterizing segments of fixed size (either time-wise or step-wise).
Auxiliary data may also become an increasingly important augmentation of relocation time series collected at relatively high frequencies in the context of compensating for observation errors associated with GPS location data. These location data currently have a CEP (circular error probable) of a few meters, depending on the particular hardware used to collect the data [9]. Averaging across $\mu$ steps however, will considerably reduce this error [53]. The addition of other data types, such as accelerometer data, are likely to increase the statistical robustness of identifying the StaMEs that underpin different types of movement, particular relatively low accelerative (e.g. gliding or striding) versus relatively high accelerative (e.g., jumping or swooping) modes of movement.
The method presented in this paper hold no matter what type of movement data are used to define a movement track, provided suitable AI clustering methods can be applied to the identification of a robust set of StaMEs. Thus, performance-measure-supported method provides a way forward to implementing a fine-scale structural analyses of animal movement tracks in the decades to come.
@@ Acknowledgments
I would like to thank Varun Sethi for constructive discussions regarding the development of the approach taken here and for taking the time to provide comments on this manuscript, as well as Richard Salter and Orr Spiegel for discussions over the past year regarding animal track segmentation. In addition, and in approximate chronological order, I have benefited over the past 20 years from the many discussions I have had with the following students and colleagues who have influenced my thinking within the field of spatial and movement ecology: Norman Owen-Smith, Jessica Redfern, Peter Baxter, Chris Wilmers, Paul Cross, Jamie Lloyd-Smith, George Wittemyer, Sadie Ryan, Shirli Bar-David, David Saltz, Ran Nathan, Craig Tambling, Leo Polansky, Andy Lyons, Wendy Turner, Steve Bellan, Alex Boettiger, Richard Salter, Werner Kiliam, Royi Zidon, Scott Fortmann-Roe, Miri Tsalyuk, Briana Abrams, Dana Seidel, Eric Dougherty, Colin Carlson, Krti Tallam, Jason Blackburn, Wayne Linklater, Yen-Hua Huang and Ludovica Luisa Vassat.